OpenStudy (anonymous):

Solve:x^4-34x^2=-225 a. no solution b.3,-5 c.3,-3,5,-5 d.3,-3

OpenStudy (anonymous):

i got x=+/- 5 and +/- 3 so the answer is c :)

OpenStudy (anonymous):

x^(4)-34x^(2)=-225 x^(4)-34x^(2)+225=0 (x^(2)-9)(x^(2)-25)=0 (x-3)(x+3)(x^(2)-25)=0 (x-3)(x+3)(x-5)(x+5)=0 x-3=0_x+3=0_x-5=0_x+5=0 x=3,-3,5,-5